The American Heart Association is about to conduct an anti-smoking campaign and wants to know the fraction of Americans over 22 who smoke.

Step 2 of 2 :  

Suppose a sample of 966 Americans over 22 is drawn. Of these people, 802 don't smoke. Using the data, construct the 98% confidence interval for the population proportion of Americans over 22 who smoke. Round your answers to three decimal places.
Answer #1

Solution :

n = 966

x = 802

\hat p = x / n = 802 / 900 = 0.830

1 -\hat p = 1 - 0.830 = 0.170

At 98% confidence level the z is ,

\alpha  = 1 - 98% = 1 - 0.98 = 0.02

\alpha / 2 = 0.02 / 2 = 0.01

Z\alpha/2 = Z0.01 = 2.326

Margin of error = E = Z\alpha / 2 * \sqrt ((\hat p * (1 - \hat p )) / n)

= 2.326 * (\sqrt((0.830 * 0.170) / 966 )

= 0.028

A 98 % confidence interval for population proportion p is ,

\hat p- E < P <\hat p + E

0.830 - 0.028 < p < 0.830 + 0.028


0.802 < p < 0.858

The 98% confidence interval for the population proportion p is : ( 0.802 , 0.858)